OpenStudy (anonymous):

A recipe for 1 lb of trail mix calls for 5.7 oz of raisins. How many ounces of raisons are needed to make 3 lb of trail mix? A. 1.9 oz B. 8.7 oz C. 17.1 oz D. 22.5 oz

OpenStudy (anonymous):

5.7 = 1lb 11.4 = 2lb ? = 3lb 5.7 * 3 = ?

OpenStudy (texaschic101):

set up a proportion 1/5.7 = 3/x --- 1 lb tm to 5.7 oz = 3 lbs of tm tp x oz cross multiply (1)(x) = (5.7)(3) x = 5.7 * 3 x = you can do the math...just multiply

OpenStudy (anonymous):

Thanks!!!! @texaschic101 @silverr

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (anonymous):

yes, no problem !